OpenStudy (anonymous):

4n - 2n = 18 = 2 ÷ 18 = n=9 (Is that right?)

OpenStudy (alexwee123):

yep the answer's correct but you work is a little off

OpenStudy (anonymous):

4n-2n =18 2n=18 n=18/2=9

OpenStudy (alexwee123):

18÷2 not 2÷18 :)

OpenStudy (anonymous):

yah that is how i got 9